# VideoGuide: Improving Video Diffusion Models without Training Through a Teacher’s Guide

### v2

@@ 4.1 Comparison results
Qualitative results for various prompts and base models are shown in Fig. 3. Samples from the base model show impairment in temporal consistency, such as fluctuation in color or abrupt change in subject appearance. FreeInit moderately solves the problem of temporal consistency but at the cost of considerable degradation in imaging quality, such as smoothing out of textural details. In contrast, the proposed VideoGuide significantly enhances temporal consistency without loss of imaging quality or motion smoothness. Furthermore, VideoGuide solves sudden frame shifts frequently observed in LaVie by providing smooth frame transitions, explained in Appendix E. Detailed explanation of base models used and additional qualitative results are included in Appendix A and Appendix E.
Method Subject consistency ($\uparrow$) Background Consistency ($\uparrow$) Imaging Quality ($\uparrow$) Motion Smoothness ($\uparrow$)
AnimateDiff (guo2023animatediff) $0.9183$ $0.9437$ $\underline{0.6647}$ $0.9547$
AnimateDiff + FreeInit (wu2023freeinit) $0.9487$ $0.9604$ $0.6173$ $0.9705$
AnimateDiff + Ours (with AnimateDiff) $\underline{0.9596}$ $\underline{0.9642}$ $0.6526$ $\underline{0.9760}$
AnimateDiff + Ours (with VideoCrafter2) $\bm{0.9614}$ $\bm{0.9664}$ $\bm{0.6671}$ $\bm{0.9772}$
LaVie (wang2023lavie) $0.9534$ $0.9599$ $0.6750$ $0.9658$
LaVie + FreeInit (wu2023freeinit) $0.9625$ $\underline{0.9643}$ $0.6533$ $\bm{0.9757}$
LaVie + Ours (with Lavie) $\underline{0.9629}$ $\bm{0.9652}$ $\underline{0.6780}$ $\underline{0.9725}$
LaVie + Ours (with VideoCrafter2) $\bm{0.9635}$ $\underline{0.9643}$ $\bm{0.6796}$ $0.9723$
@@
In quantitative comparison, our method demonstrates superior performance over the base model, achieving improvements in both subject and background consistency. When using AnimateDiff as the base model, our approach shows best results for all key metrics. There is a notable enhancement in temporal consistency compared to baselines, and such increase is not at the cost of imaging quality or motion smoothness. Our method is shown to actually improve both factors when VideoCrafter2 is used as the guiding model. A small decrease in imaging quality can be observed for the self-guided case, but the difference is minimal compared to the notable decrease in imaging quality for FreeInit. When using LaVie as the base model, our approach still shows a reliable increase in subject and background consistency. Note that increase is relatively smaller due to a higher base consistency. Furthermore, our method successfully maintains imaging quality and improves motion smoothness. Such results conform with our original purpose to create a method for improving temporal consistency while preserving imaging quality and motion smoothness. Additionally we conduct a user study to prove the effectiveness of our approach regarding Text Alignment, Overall Quality, and Smooth And Dynamic Motion, further explained in Appendix C.
@@
FreeInit $51.88$ $28.18$
Ours (self-guided) $\bm{21.02}$ $\bm{8.99}$
Ours (VC-guided) $\underline{29.22}$ $\underline{13.43}$
Table 2: Inference time for video generation($s$). Both the self-guided case and the VideoCrafter2-guided case show significant decrease in inference time. Bold: best, underline: second best.
@@ 5.1 Ablation Study
Importance of Guidance Scale $w$. Recent study (chung2024cfg++) demonstrates that employing a high CFG scale ($w>1.0$) in the early timesteps of diffusion sampling leads to off-manifold behavior. This phenomenon results in denoised samples exhibiting problems such as color saturation and abrupt transitions, which negatively affect the interpolation between samples during these timesteps. We solve this by applying a lower guidance scale $w$ during the early stages of sampling, ensuring smoother interpolation between the denoised samples. As illustrated in Fig. 4 (a), when using a high CFG scale ($w=7.5$), the influence of the guiding diffusion model becomes minimal due to significant color saturation, making it difficult for the output of the guiding model to be reflected effectively. In contrast, as illustrated in Fig. 4 (b), a lower CFG scale ($w=0.8$) facilitates smoother interpolation between the sampling diffusion model and the guiding diffusion model. This highlights the importance of clean interpolation in our method, as improper guidance can lead to sub-optimal performance. Further analysis about CFG and CFG++ can be found in Appendix B.
@@
Parameter Selection. An analysis is performed to assess how varying parameters of the guiding diffusion model impacts temporal consistency. Specifically, we examine the effects of three factors: interpolation scale $\beta$, number of interpolation steps $I$, and number of guidance sampling steps $\tau$. Temporal consistency is evaluated for both Subject Consistency (SC) and Background Consistency (BC). To secure efficient sampling time, we limit maximum values to $\tau=10$ and $I=5$.
Our ablation studies prove that all three parameters are closely related to temporal consistency. Decrease in interpolation scale $\beta$, which is analogous to increase in the influence of the guiding diffusion model, leads to improved subject and background consistency. Note that the minimum value of $\beta$ is constrained to $0.5$ to mitigate the risk of distribution shift. Increasing the number of interpolation steps $I$ also leads to improvement in temporal consistency, which proves that our interpolation scheme is indeed effective. Furthermore, increasing the number of guidance sampling steps $\tau$ enhances consistency, indicating that blending intermediate latents with better-denoised versions enhances overall consistency as expected (i.e., ${\boldsymbol{z}}_{0|t-\tau}\approx{\boldsymbol{z}}_{r}$). Such ablation study highlights the trade-off between consistency improvement and computational efficiency, offering insight into optimal parameter settings for the guiding diffusion model.
@@ 5.2 Prior Distillation
Degraded performance due to a substandard data prior is an issue only solvable through extra training. However VideoGuide provides a workaround to this matter by enabling the utilization of a superior data prior. Fig. 5 demonstrates example cases. For all instances, generated samples are guided towards a result of better text coherence while maintaining the style of the original data domain. Additional examples of prior distillation are provided in Appendix E.
@@ 6 Conclusion
Appendix A Experimental Details
A.1 Prompt Selection
@@
In all experiments, we utilize 800 prompts from various categories in VBench (huang2023vbench) to evaluate the model’s ability to generate across diverse categories.
A.2 Hyperparameter Selection
@@
We employ a classifier-free guidance (CFG) scale of $7.5$ during inference for both base models (AnimateDiff, LaVie) and FreeInit-applied cases. During interpolation of the denoised samples, we apply CFG++ reverse sampling with a guidance scale of $w=0.8$ in DDIM $50$-step sampling. After completing the interpolation step, we revert to CFG reverse sampling with a CFG scale of $7.5$. In FreeInit, we use a Butterworth filter with a normalized frequency of $0.25$, filter order $n=4$, and perform $5$ iterations, as recommended in prior work. The same filter is applied in our experiments with FreeInit. For AnimateDiff, we configure the guiding model with parameters $I=5$, $\beta=0.5$, and $\tau=10$. In the case of LaVie, we set $I=3$, $\beta=0.5$, and $\tau=10$ to optimize inference speed. Additionally, the $\tau$ intervals are not uniformly spaced as in the standard $50$-step DDIM sampling. To better leverage temporally consistent samples, we divide the remaining interval into $25$ steps for reverse sampling during guidance steps. Also, we found that applying renoising to guidance sampling is more effective in improving consistency in the case of self-guidance. Therefore, we incorporated renoising during self-guidance in a similar manner as when using a external model for guidance.
A.3 Figure Explanation
@@
(c) AnimateDiff with pretrained T2I model ToonYou.
(d) AnimateDiff with pretrained T2I model FilmVelvia.
(e) LaVie.
(f) LaVie.
Base model used for Figure 5: AnimateDiff with pretrained T2I model ToonYou.
Appendix B Quantitative Analysis of CFG and CFG++
Metrics Base FreeInit Ours
CFG CFG++ CFG CFG++ CFG Interp. CFG++ Interp.
Subject Consistency ($\uparrow$) $0.9183$ $0.9176$ $0.9487$ $0.9473$ $0.9598$ $0.9614$
Background Consistency ($\uparrow$) $0.9437$ $0.9435$ $0.9604$ $0.9604$ $0.9635$ $0.9664$
Table 4: Comparison of consistency metrics between CFG and CFG++ in AnimateDiff. Results indicate that interpolating denoised samples with CFG++ has a larger impact on improving both subject and background consistency.
There may be concerns that the effectiveness of our method in improving consistency stems from the use of the CFG++ algorithm itself. To address this, we provide results for using CFG and CFG++ across the Base Model, Base Model + FreeInit, and Base Model + VideoGuide. The results demonstrate that CFG++ is particularly effective for interpolation. As shown in Tab. 4, metrics for Base and FreeInit decrease when CFG++ is used, and metrics improve only when CFG++ is applied to our interpolation scheme. This implies the significant positive impact on consistency of CFG++ within the proposed interpolation scheme, especially compared to CFG. Also, this supports the idea, as discussed earlier in Sec. 5.1, that smooth interpolation of denoised samples positively impacts model performance.
Appendix C User Study
We conduct a user study to evaluate generated video samples using three criteria: Text Alignment, Overall Quality, and Smooth And Dynamic Motion, with all metrics scored on a 1 to 5 scale. A total of 30 participants provided ratings for each metric, offering comprehensive feedback on the generated videos.

### v3

@@ 4.1 Comparison Results
Qualitative results for various prompts and base models are shown in Fig. 3. Samples from the base model show impairment in temporal consistency, such as fluctuation in color or abrupt change in subject appearance. FreeInit [wu2023freeinit] moderately solves the problem of temporal consistency but at the cost of considerable degradation in imaging quality, such as smoothing out of textural details. In contrast, the proposed VideoGuide significantly enhances temporal consistency without loss of imaging quality. Furthermore, VideoGuide solves sudden frame shifts frequently observed in LaVie by providing smooth frame transitions. Additional qualitative results are included in Supplementary Material.
Method Subject consistency ($\uparrow$) Background Consistency ($\uparrow$) Imaging Quality ($\uparrow$) Motion Smoothness ($\uparrow$) Dynamic Degree ($\uparrow$)
AnimateDiff [guo2023animatediff] $0.9183$ $0.9437$ $\underline{0.6647}$ $0.9547$ $\bm{26.67}$
AnimateDiff + FreeInit [wu2023freeinit] $0.9487$ $\underline{0.9604}$ $0.6173$ $0.9705$ $\underline{19.28}$
AnimateDiff + Ours (with AnimateDiff) $\underline{0.9520}$ $0.9600$ $0.6566$ $\underline{0.9731}$ $15.25$
AnimateDiff + Ours (with VideoCrafter2) $\bm{0.9614}$ $\bm{0.9664}$ $\bm{0.6671}$ $\bm{0.9772}$ $16.78$
LaVie [wang2023lavie] $0.9534$ $0.9599$ $0.6750$ $0.9658$ $\bm{14.69}$
LaVie + FreeInit [wu2023freeinit] $0.9625$ $\underline{0.9643}$ $0.6533$ $\bm{0.9757}$ $10.69$
LaVie + Ours (with Lavie) $\underline{0.9629}$ $\bm{0.9652}$ $\underline{0.6780}$ $\underline{0.9725}$ $12.39$
LaVie + Ours (with VideoCrafter2) $\bm{0.9635}$ $\underline{0.9643}$ $\bm{0.6796}$ $0.9723$ $\underline{12.63}$
@@
Figure 3: Qualitative Comparison. VideoGuide is applied on various base models for different text prompts. For each prompt, frames of generated samples from four different models are displayed: (i) Base model (first row); (ii) Base model with FreeInit (second row); (iii) Base model with VideoGuide (self-guided case) (third row); (iv) Base model with VideoGuide (external model-guided case) (fourth row). AD, VC, LV indicate guidance models of AnimateDiff, VideoCrafter-2.0, LaVie, respectively. Samples for the base model show substandard temporal consistency, especially regarding color fluctuation and subject appearance change. Applying FreeInit improves consistency but introduces degradation in imaging quality, such as smoothing out of textural details. In contrast, applying VideoGuide significantly enhances temporal consistency while preserving imaging quality, both for the self-guided and the external model-guided case.
Quantitative comparison shows that our approach consistently outperforms the baseline model, achieving substantial improvements in both subject and background consistency. When using AnimateDiff [guo2023animatediff] as the baseline, our method delivers the best results across all key metrics, except for dynamic degree, due to an inherent trade-off. Our method greatly enhances temporal consistency with a comparatively small reduction in dynamic degree, ensuring stable temporal coherence without compromising image quality. For LaVie [wang2023lavie], our method achieves higher temporal consistency and an increased dynamic degree compared to previous methods, highlighting its ability to enhance temporal stability with minimal impact on dynamic motion.
Incorporating an external model [chen2024videocrafter2] for guidance enhances performance by achieving higher temporal consistency without compromising dynamic motion, compared to self-guided cases in both AnimateDiff and LaVie. This result indicates that using a higher-performing external model for guidance can lead to superior video quality. Additionally, our method successfully addresses challenges such as sudden visual shifts, as discussed in Supplementary Material.
@@
Baseline $\bm{11.38}$ $\bm{6.86}$
FreeInit $51.98$ $30.18$
Ours (self-guided) $\underline{21.68}$ $\underline{10.01}$
Ours (VC-guided) $29.73$ $14.07$
Table 2: Inference time for video generation($s$). Both the self-guided case and the VideoCrafter2-guided case show significant decrease in inference time compared with previous method [wu2023freeinit]. Bold: best, underline: second best.
@@
Our approach achieves these advancements without sacrificing image quality, highlighting its importance for both personalization-focused applications and high-quality video generation. Overall, VideoGuide proves to be essential for optimizing temporal coherence and enhancing overall quality across various model frameworks.
@@ 5.1 Ablation Study
Parameter Selection. An analysis is performed to assess how varying parameters of the guiding diffusion model impacts temporal consistency. Specifically, we examine the effects of three factors: interpolation scale $\beta$, number of interpolation steps $I$, number of guidance sampling steps $\tau$.
Our ablation studies prove that all three parameters are closely related to temporal consistency. Decrease in interpolation scale $\beta$ leads to improved subject and background consistency. Note that the minimum value of $\beta$ is constrained to $0.5$ to mitigate the risk of distribution shift. Increasing the number of interpolation steps $I$ also leads to improvement in temporal consistency, which proves that our interpolation scheme is indeed effective. Furthermore, increasing the number of guidance sampling steps $\tau$ enhances consistency, indicating that blending intermediate latents with better-denoised versions enhances overall consistency as expected. Such ablation study highlights the trade-off between consistency improvement and computational efficiency, offering insight into optimal parameter settings for the guiding diffusion model.
@@ 5.2 Prior Distillation
Degraded performance due to a substandard data prior is an issue only solvable through extra training. However VideoGuide provides a workaround to this matter by enabling the utilization of a superior data prior. Fig. 4 demonstrates example cases. For all instances, generated samples are guided towards a result of better text coherence while maintaining the style of the original data domain. Additional examples of prior distillation are provided in Supplemental Material.
@@ 6 Conclusion
Supplementary Material
The supplementary material is organized as follows:
  • Section A: Importance of the low-pass filter.
  • Section B: About classifier-free guidance.
  • Section C: Pseudocodes for our algorithm.
  • Section D: More experimental details.
  • Section E: More quantitative results: user study.
  • Section F: Application on DiT-based models.
  • Section G: Comparison with orthogonal methods.
  • Section H: Limitations.
  • Section I: More qualitative examples.
Appendix A Importance of Low-Pass Filter
To evaluate the role of the low-pass filter in our methodology, we conduct experiments by varying the interpolation step $I$, both with and without the low-pass filter. These experiments are averaged across 800 prompts from the VBench categories for consistent evaluation. We apply the low-pass filter for the initial 5 timesteps, based on the observation that the mid-to-late timesteps in the diffusion process focus on generating mid- and high-frequency details. Replacing these frequencies with random components via the low-pass filter in the mid-to-late timesteps would result in degraded video quality, necessitating the early-timestep limitation. All corresponding results are presented in Fig. 5.
We measure the effects of the filter on Subject Consistency, Background Consistency, and Imaging Quality. Both Subject Consistency and Background Consistency steadily improves as the number of interpolation steps increases, demonstrating the effectiveness of our interpolation scheme in enhancing temporal coherence. Meanwhile, Imaging Quality is maintained up to approximately 10 interpolation steps without the low-pass filter. Beyond this point, a significant drop in quality is observed, indicating that excessive interpolation exacerbates the blurring effects caused by prolonged SDS optimization, as noted earlier in this work.
The improvement in consistency is significantly accelerated when using the low-pass filter. This acceleration is achieved while mitigating the decline in imaging quality typically associated with increased interpolation steps. Furthermore, application of the filter also reduces computational overhead during interpolation. Specifically, the consistency achieved at $I=4$ with the filter is comparable to the consistency achieved at $I=50$ without the filter, offering approximately a 7-fold reduction in inference time. Such results prove the effectiveness of the low-pass filter in balancing consistency improvement, imaging quality preservation, and computational efficiency.
Figure 5: Comparison of Subject Consistency, Background Consistency, and Imaging Quality across interpolation steps ($I$) with and without the application of the low-frequency filter. Results indicate that the low-frequency filter accelerates convergence towards consistency while maintaining imaging quality.
Appendix B Classifier-Free Guidance
Off-Manifold Behavior of CFG Recent study [chung2024cfg++] demonstrates that employing a high Classifier- Free Guidance (CFG) [ho2021classifierfree] scale ($w>1.0$) in the early timesteps of diffusion sampling leads to off-manifold behavior. This phenomenon results in denoised samples exhibiting problems such as color saturation and abrupt transitions, which negatively affect the interpolation between samples during these timesteps. We solve this by applying a lower guidance scale $w$ during the early stages of sampling, ensuring smoother interpolation between the denoised samples. As illustrated in Fig. 6 (a), when using a high CFG scale ($w=7.5$), the influence of the guiding diffusion model becomes minimal due to significant color saturation, making it difficult for the output of the guiding model to be reflected effectively. In contrast, as illustrated in Fig. 6 (b), a lower CFG scale ($w=0.8$) facilitates smoother interpolation between the sampling diffusion model and the guiding diffusion model.
Configuration SC ($\uparrow$) BC ($\uparrow$)
Base
      + CFG $0.9183$ $0.9437$
      + CFG++ $0.9176$ $0.9435$
FreeInit
      + CFG $0.9487$ $0.9604$
      + CFG++ $0.9473$ $0.9604$
Ours
      + CFG Interp. $0.9598$ $0.9635$
      + CFG++ Interp. $0.9614$ $0.9664$
Table 4: Comparison of consistency metrics (SC: Subject Consistency, BC: Background Consistency) across different configurations using CFG and CFG++ in AnimateDiff. Our approach with interpolated CFG++ achieves the best performance, significantly enhancing both subject and background consistency.
We provide quantitative analysis for using CFG and CFG++ across the Base Model, Base Model + FreeInit, and Base Model + VideoGuide (Ours) during the interpolation. As shown in Tab. 4, metrics for Base and FreeInit decrease when CFG++ is used, and metrics improve only when CFG++ is applied to our interpolation scheme. This implies the significant positive impact on consistency of CFG++ within the proposed interpolation scheme, especially compared to CFG. Also, this supports the idea that smooth interpolation of denoised samples positively impacts model performance, as discussed above.
Appendix C Pseudo Code
Pseudo codes regarding our algorithm are provided. For clarity, the pseudo code describing our algorithm adopts the CFG++ reverse sampling framework for the entire process.
Algorithm 1 VideoGuide with Sampling Diffusion Model
1:   Initialize ${\bm{z}}_{T}\sim\mathcal{N}(0,{\mathbf{I}})$
2:   for $t=T,\dots,1$ do
3:      $\hat{{\bm{\epsilon}}}_{\theta}({\bm{z}}_{t},t)={\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t,\phi)+\lambda[{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t,c)-{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t,\phi)]$
4:      ${\bm{z}}_{0|t}=({\bm{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{{\bm{\epsilon}}_{\theta}}({\bm{z}}_{t},t))/\sqrt{\bar{\alpha}_{t}}$
5:      ${\bm{z}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},\quad\textit{where}\quad{\bm{\epsilon}}\sim N(0,{\mathbf{I}})$
6:      if $T-t<I$ then
7:         for $j=0,\dots,\tau$ do
8:            ${\bm{z}}_{t-j-1}=\sqrt{\bar{\alpha}_{t-j-1}}{\bm{z}}_{0|t-j}+\sqrt{1-\bar{\alpha}_{t-j-1}}{\bm{\epsilon}}_{\theta}({\bm{z}}_{t-j},t-j,\phi)$
9:         end for
10:         ${\bm{z}}^{\prime}_{0|t}=\beta\cdot{\bm{z}}_{0|t}+(1-\beta)\cdot{\bm{z}}_{0|t-\tau}$
11:         ${\bm{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}^{\prime}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}}{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t,\phi)$
12:         ${\bm{z}}_{t-1}=\textit{LPF}_{\gamma}({\bm{z}}_{t-1})+\textit{HPF}_{\gamma}({\bm{\epsilon}}),\quad\textit{where}\quad{\bm{\epsilon}}\sim N(0,{\mathbf{I}})$
13:      else
14:         ${\bm{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}}{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t,\phi)$
15:      end if
16:   end for
17:   Output: Final video ${\bm{z}}_{0}$
Algorithm 2 VideoGuide with Guiding Diffusion Model
1:   Initialize ${\bm{z}}_{T}\sim\mathcal{N}(0,{\mathbf{I}})$
2:   for $t=T,\dots,1$ do
3:      $\hat{{\bm{\epsilon}}}_{\theta}({\bm{z}}_{t},t)={\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t,\phi)+\lambda[{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t,c)-{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t,\phi)]$
4:      ${\bm{z}}_{0|t}=({\bm{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{{\bm{\epsilon}}_{\theta}}({\bm{z}}_{t},t))/\sqrt{\bar{\alpha}_{t}}$
5:      ${\bm{z}}^{(G)}_{t}=\sqrt{\bar{\alpha}^{(G)}_{t}}z_{0|t}+\sqrt{1-\bar{\alpha}^{(G)}_{t}}{\bm{\epsilon}},\quad\textit{where}\quad{\bm{\epsilon}}\sim N(0,{\mathbf{I}})$
6:      if $T-t<I$ then
7:         for $j=0,\dots,\tau$ do
8:            ${\bm{z}}_{0|t-j}^{(G)}=({\bm{z}}_{t-j}^{(G)}-\sqrt{1-\bar{\alpha}_{t-j}^{(G)}}\hat{{\bm{\epsilon}}}_{\psi}({\bm{z}}_{t-j}^{(G)},t-j)/\sqrt{\bar{\alpha}_{t-j}^{(G)}}$
9:            ${\bm{z}}_{t-j-1}^{(G)}=\sqrt{\bar{\alpha}_{t-j-1}^{(G)}}{\bm{z}}_{0|t-j}^{(G)}+\sqrt{1-\bar{\alpha}_{t-j-1}^{(G)}}{\bm{\epsilon}}_{\psi}({\bm{z}}_{t-j}^{(G)},t-j,\phi)$
10:         end for
11:         ${\bm{z}}^{\prime}_{0|t}=\beta\cdot{\bm{z}}_{0|t}+(1-\beta)\cdot{\bm{z}}_{0|t-\tau}^{(G)}$
12:         ${\bm{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}^{\prime}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}}{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t,\phi)$
13:         ${\bm{z}}_{t-1}=\textit{LPF}_{\gamma}({\bm{z}}_{t-1})+\textit{HPF}_{\gamma}({\bm{\epsilon}}),\quad where\quad{\bm{\epsilon}}\sim N(0,{\mathbf{I}})$
14:      else
15:         ${\bm{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}}{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t,\phi)$
16:      end if
17:   end for
18:   Output: Final video ${\bm{z}}_{0}$
Appendix D Experimental Details
D.1 Prompt Selection
@@
In all experiments, we utilize 800 prompts from various categories in VBench [huang2023vbench] to evaluate the model’s ability to generate across diverse categories.
D.2 Hyperparameter Selection
@@
We employ a classifier-free guidance (CFG) scale of $7.5$ during inference for both base models (AnimateDiff, LaVie) and FreeInit-applied cases. During interpolation of the denoised samples, we apply CFG++ reverse sampling with a guidance scale of $w=0.8$ in DDIM $50$-step sampling. After completing the interpolation step, we revert to CFG reverse sampling with a CFG scale of $7.5$. In FreeInit, we use a Butterworth filter with a normalized frequency of $0.25$, filter order $n=4$, and perform $5$ iterations, as recommended in prior work. The same filter is applied in our experiments with FreeInit. For AnimateDiff, we configure the guiding model with parameters $I=5$, $\beta=0.5$, and $\tau=10$. In the case of LaVie, we set $I=3$, $\beta=0.5$, and $\tau=10$ to optimize inference speed. Additionally, the $\tau$ intervals are not uniformly spaced as in the standard $50$-step DDIM sampling. To better leverage temporally consistent samples, we divide the remaining interval into $25$ steps for reverse sampling during guidance steps.
D.3 Figure Explanation
@@
(c) LaVie.
(d) LaVie.
Base model used for Figure 4:
AnimateDiff with pretrained T2I model ToonYou.
Appendix E User Study
We conduct a user study to evaluate generated video samples using three criteria: Text Alignment, Overall Quality, and Smooth and Dynamic Motion, with all metrics scored on a 1 to 5 scale. A total of 30 participants provided ratings for each metric, offering comprehensive feedback on the generated videos. Tab. 5 shows that our method surpasses the baseline and previous work in all evaluated aspects.